\begin{thebibliography}{00}

\bibitem{Beirao da Veiga_95}
     H. Beir\~ao da Veiga,
     A new regularity class for the Navier-Stokes equations in $\bbR^n$,
     \emph{Chinese Ann. Math. Ser. B,}
     {\bf 16} (1995), 407--412.

\bibitem{Beirao da Veiga_02}
     H. Beir\~ao da Veiga, L.C. Berselli,
     On the regularizing effect of the vorticity direction in incompressible viscous flows,
     \emph{Differential Integral Equations,}
     {\bf 15} (2002), 345--356.

\bibitem{CT10_p3u3}
    C.S. Cao, E.S. Titi,
    Global regularity criterion for the 3D Navier-Stokes equations involving one entry of the velocity gradient tensor,
    \emph{Arch. Rational Mech. Anal.,}
    {\bf 202} (2011), 919--932.

\bibitem{CL01_p}
    D. Chae, J. Lee,
    Regularity criterion in terms of pressure for the Navier-Stokes equations,
    \emph{ Nonlinear Anal.,}
    {\bf 46} (2001), 727--735.

\bibitem{Constantin_Fefferman_93}
     P. Constantin, C. Fefferman,
     Direction of vorticity and the problem of global regularity for the Navier-Stokes equations,
     \emph{Indiana Univ. Math. J.,}
     {\bf42} (1993), 775--789.

\bibitem{Escauriaza_Seregin_Sverak_03}
     L. Escauriaza, G. Seregin, V. \v Sver\'ak,
     Backward uniqueness for parabolic equations,
     \emph{Arch. Ration. Mech. Anal.,}
     {\bf169} (2003), 147--157.

\bibitem{FJN08_p}
    J.S. Fan, S. Jiang,  G.X. Ni,
    On regularity criteria for the $n$-dimensional Navier-Stokes equations in terms of the pressure,
    \emph{J. Differential Equations,}
    {\bf 244} (2008), 2963--2979.

\bibitem{He_Gala_11_NARWA}
    X.W. He, S. Gala,
    Regularity criterion for the weak solutions to the Navier-Stokes equations in terms of the pressure in the class $L^2(0,T;\dot B^{-1}_{\infty,\infty}(\bbR^3))$,
    {\em Nonlinear Analysis: RWA.},
    {\bf 12} (2011), 3602--3607.

\bibitem{Hopf}
    E. Hopf,
    \"Uber die Anfangswertaufgabe f\"ur die hydrodynamischen Grundgleichungen,
    \emph{ Math. Nachr.},
    {\bf 4} (1951), 213--231.

\bibitem{Leray}
    J. Leray,
    Sur le mouvement d'un liquide visqueux emplissant l'espace,
    \emph{ Acta Math.},
    {\bf 63} (1934), 193--248.

\bibitem{MGO_97}
    Y. Meyer, P. Gerard, F. Oru,
    In\'egalit\'es de Sobolev pr\'ecis\'ees, S\'eminaire \'E quations aux d\'eriv\'ees partielles (Polytechnique)
    (1996--1997), Exp. No. 4, 8 pp.

\bibitem{Neustupa_Penel}
    J. Neustupa, P. Penel,
    Anisotropic and geometric criteria for interior regularity of weak solutions to the $3$D Navier¨CStokes equations,
    in Mathematical Fluid Mechanics (Recent Results and Open Problems),
    Advances in Mathematical Fluid Mechanics,
    edited by J. Neustupa, and P. Penel (Birkh\"auser, Basel-Boston-Berlin,
    (2001), 239--267.

\bibitem{Penel_Pokorny}
    P. Penel, M. Pokorn\'y,
    Some new regularity criteria for the Navier-Stokes equations containing the gradient of velocity,
    \emph{Appl. Math.,}
    {\bf49} (2004), 483--493.

\bibitem{Prodi_59}
     G. Prodi,
     Un teorema di unicit\'a per le equazioni di Navier-Stokes,
     \emph{Ann. Mat. Pura Appl.,}
     {\bf 48} (1959), 173--182.

\bibitem{Serrin_62}
    J. Serrin,
    On the interior regularity of weak solutions of the Navier-Stokes equations.
    \emph{Arch. Rational Mech. Anal.,}
    {\bf 9} (1962), 187--191.

\bibitem{Serrin_63}
    J. Serrin,
    The initial value problems for the Navier-Stokes equations.
    in Nonlinear Problems, edited by R. E. Langer (University of Wisconsin Press, Madison, WI, 1963.

\bibitem{Zhang_Chen_05}
    Z.F. Zhang and Q.L. Chen,
    Regularity criterion via two components of vorticity on weak solutions to the Navier-Stokes equations in $\bbR^3$,
    \emph{J. Differential Equations,}
    {\bf 216} (2005), 470--481.

\bibitem{Zhang_arXiv_Hess}
    Z.J. Zhang,
    A regularity criterion for the Navier-Stokes equations via two entries of the velocity Hessian tensor,
    arXiv: 1103.1196.

\bibitem{Zhang_arXiv_u3}
    Z.J. Zhang,
    A Serrin-type regularity criterion for the Navier-Stokes equations via one velocity component,
    Commun. Pure Appl. Anal., {\bf 12} (2013), 117--124.


\bibitem{ZhYLGL}
    Z.J. Zhang, Z.A. Yao, P. Li, C.C. Guo, M. Lu,
    Two new regularity criteria for the 3D Navier-Stokes equations via two entries of the velocity gradient tensor,
    \emph{Acta Appl. Math.,}
    (2012), doi: 10.1007/s10440-012-9712-4.

\bibitem{Zhou_05_one component}
    Y. Zhou,
    A new regularity criterion for weak solutions to the Navier-Stokes equations,
    {\em  J. Math. Pures Appl.,}
    {\bf  84} (2005), no. 11, 1496--1514.

\bibitem{ZP09_pu3}
    Y. Zhou, M. Pokorn\'y,
    On a regularity criterion for the Navier-Stokes equations involving gradient of one velocity component,
    \emph{ J. Math. Phys.,}
    {\bf 50}
    (2009), 123514, 11 pp.

\bibitem{ZP10_u3_p3u3}
    Y. Zhou, M. Pokorn\'y,
    On the regularity of the solutions of the Navier-Stokes equations via one velocity component,
    \emph{ Nonlinearity,}
    {\bf 23}
    (2010), 1097--1107.

%\bibitem{Fan_Zhou_09_AML}
%    J.S. Fan, Y. Zhou,
%    A note on the regularity criterion for the $3$D Boussinesq equations with partial viscosity,
%    {\em Appl. Math. Lett.},
%    {\bf 22} (2009), 802--805.
%\bibitem{Ish_Mor_99_MMAS}
%    N. Ishihara, H. Morimoto,
%    Remarks on the blow-up criterion for the $3$D Boussinesq equations,
%    {\em Math. Models Methods Appl.},
%    {\bf 9} (1999), 1323--1332.
%
%\bibitem{Lem02}
%    P.G. Lemari\'e-Rieusset,
%    Recent developments in the Navier-Stokes problem,
%    Chapman \& Hall/CRC, Boca Raton, FL, 2002.
%
%
%\bibitem{NNP}
%    J. Neustupa,
%    A. Novotn\'y,
%    P. Penel,
%    An interior regularity of a weak solution to the Navier-Stokes equations in dependence on one component of velocity,
%    Topics in mathematical fluid mechanics,
%    {\em  Quad. Mat.},
%    {\bf 10} (2002), 163--183.
%
%\bibitem{QDY_11_CNSNS}
%    H. Qiu, Y. Du, Z.A. Yao,
%    Blow-up criteria for the $3$D Boussinesq equations in the multiplier spaces,
%    {\em Commun. Nonlinear Sci. Numer. Simul.},
%    {\bf 16}
%    (2011), 1820--1824.
%
%\bibitem{QDY_10_AA}
%    H. Qiu, Y. Du, Z.A. Yao,
%    Serrin-type blow-up criteria for three-dimensional Boussinesq equations,
%    {\em Appl. Anal.},
%    {\bf 89} (2010), 1603--1613.
%
%
\end{thebibliography}

\documentclass[12pt]{elsarticle}

\usepackage{amsmath,amsthm,amssymb,dsfont,a0size,bm,indentfirst,wasysym}
\usepackage{stmaryrd,CJK,color,graphicx}
\usepackage[toc,page,title,titletoc,header]{appendix}
\usepackage{paralist}
\usepackage{epsfig}
\usepackage[colorlinks=true]{hyperref}
\hypersetup{citecolor=red}

\usepackage{remreset}

\allowdisplaybreaks

\newcommand{\bex}{\begin{eqnarray*}}
\newcommand{\eex}{\end{eqnarray*}}
\newcommand{\be}{\begin{eqnarray}}
\newcommand{\ee}{\end{eqnarray}}

\newcommand{\beex}{\begin{equation*}}
\newcommand{\eeex}{\end{equation*}}

\newcommand{\bee}{\begin{equation}}
\newcommand{\eee}{\end{equation}}
\newcommand{\bea}{\begin{aligned}}
\newcommand{\eea}{\end{aligned}}

\newcommand{\ba}{\begin{array}}
\newcommand{\ea}{\end{array}}
\newcommand{\bi}{\begin{itemize}}
\newcommand{\ei}{\end{itemize}}
\newcommand{\bn}{\begin{enumerate}}
\newcommand{\en}{\end{enumerate}}
\newcommand{\sex}[1]{\left(#1\right)}

\newcommand{\sed}[1]{\left\{#1\right\}}

\newcommand{\sen}[1]{\left\Vert#1\right\Vert}

\renewcommand{\theequation}{\arabic{equation}}

\newtheorem{thm}{\indent Theorem}

\newtheorem{defn}[thm]{\indent Definition}

\newcommand{\bp}{\begin{proof}}
\newcommand{\ep}{\end{proof}}

\newcommand{\ve}{\varepsilon}

\def\n{\nabla}

\def\lap{\triangle}

\def\calF{\mathcal{F}}

\def\calS{\mathcal{S}}

\def\calZ{\mathcal{Z}}

\def\bbu{\bm{u}}

\def\bbN{\mathbb{N}}

\def\bbR{\mathbb{R}}

\def\bbZ{\mathbb{Z}}

\def\rd{\mathrm{\,d}}

\allowdisplaybreaks
\def\ve{\varepsilon}

\makeatletter
\def\@eqnnum{{\normalfont \color{red} (\theequation)}}
\makeatother

\makeindex

\begin{document}

\begin{frontmatter}

\title{A smallness regularity criterion for the $3$D Navier-Stokes equations in the largest class}

\author[a]{Zujin Zhang\fnref{z}\corref{cor}}

\address[a]{
School of Mathematics and Computer Science,
Gannan Normal University\\
Ganzhou 341000, P.R. China}

\cortext[cor]{Corresponding author}

\fntext[z]{zhangzujin361@163.com}

\begin{abstract}
In this paper, we consider the three-dimensional Navier-Stokes equations, and show
that if the $\dot B^{-1}_{\infty,\infty}$-norm of the velocity field is sufficiently
small, then the solution is in fact classical.
\end{abstract}

\begin{keyword}
Navier-Stokes equations\sep
regularity criterion\sep
Besov spaces

\MSC[2010]
35B65\sep
76B03\sep
76D03

\end{keyword}

\end{frontmatter}

\section{Introduction}
\label{sect:intro}
    Consider the following three-dimensional ($3$D) Navier-Stokes equations:
    \bee
    \label{NSE}
    \left\{
    \ba{l}
    \bbu_t+(\bbu\cdot\n)\bbu-\lap \bbu+\n \pi=0,\\
    \n\cdot \bbu=0,\\
    \bbu(x,0)=\bbu_0,
    \ea\right.
    \eee
    where $\bbu=(u_1(x,t),u_2(x,t),u_3(x,t))$ is the fluid velocity,
    $\pi=\pi(x,t)$ is a scalar pressure; and $\bbu_0$ is the prescribed
    initial velocity filed satisfying the compatibility condition $\n\cdot\bbu_0=0$.

    The existence of a global weak solution
    \bex
    \bbu\in L^\infty(0,T;L^2(\bbR^3))
    \cap
    L^2(0,T;H^1(\bbR^3))
    \eex
    to \eqref{NSE} has long been established by Leray
    \cite{Leray}, see also Hopf \cite{Hopf}.
    But the issue of regularity and uniqueness of $\bbu$ remains open.
    Initialed by Serrin \cite{Serrin_62,Serrin_63} and
    Prodi \cite{Prodi_59}, there have been a
    lot of literatures devoted to finding sufficient conditions
    to ensure $\bbu$ to be smooth, see, e.g., \cite{Beirao da Veiga_95,Beirao da Veiga_02,CT10_p3u3,CL01_p,
    Constantin_Fefferman_93,Escauriaza_Seregin_Sverak_03,FJN08_p,
    He_Gala_11_NARWA,Neustupa_Penel,Penel_Pokorny,Zhang_Chen_05,Zhang_arXiv_Hess,Zhang_arXiv_u3,
    ZhYLGL,Zhou_05_one component,ZP09_pu3,ZP10_u3_p3u3} and references cited therein. Noticeably,
    the following Ladyzhenskaya-Prodi-Serrin condition
    (\cite{Escauriaza_Seregin_Sverak_03,Prodi_59,Serrin_62,Serrin_63}):
    \be\label{reg_LSP}
    \bbu\in L^p(0,T;L^q(\bbR^3)),\mbox{ with }\frac{2}{p}+\frac{3}{q}=1,\quad
    3\leq q\leq \infty
    \ee
    can ensure the smoothness of the solution.

    Note that the limiting case $L^\infty(0,T;L^3(\bbR^3))$ in \eqref{reg_LSP}
    does not fall into the framework of standard
    energy method, which was proved by Escauriaza, Seregin and
    \v Sver\'ak \cite{Escauriaza_Seregin_Sverak_03} using backward uniqueness theorem.
    Due to the fact that
    \bex
    L^3(\bbR^3)\subset \dot B^{-1}_{\infty,\infty}(\bbR^3),\mbox{ but }
    L^3(\bbR^3)\neq \dot B^{-1}_{\infty,\infty}(\bbR^3),
    \eex
    we shall consider in this paper the regularity of solutions of
    \eqref{NSE} in $\dot B^{-1}_{\infty,\infty}(\bbR^3)$.
    However, we could not prove a regularity criterion as $L^\infty(0,T;\dot B^{-1}_{\infty,\infty}(\bbR^3))$,
    since the function in $\dot B^{-1}_{\infty,\infty}(\bbR^3)$ has no decay at infinity, which ensures that the
    solution is smooth outside an big ball centered at origin so that the backward uniqueness theorem can be applied.

    Before we state the precise result, let us recall the weak formulation of \eqref{NSE}.

    \begin{defn}\label{defn:wf}
    Let $\bbu_0\in L^2(\bbR^3)$ satisfying $\n\cdot\bbu_0=0$,
    $T>0$. A measurable vector-valued function $\bbu$
    defined in $[0,T]\times\bbR^3$ is said to be a weak solution to \eqref{NSE} if
    \bn
    \item $\bbu\in L^\infty(0,T;L^2(\bbR^3))\cap L^2(0,T;H^1(\bbR^3))$;
    \item $\bbu$ satisfies $\eqref{NSE}_{1,2}$ in the sense of distributions;
    \item $\bbu$ satisfies the energy inequality:
        \bex
        \sen{\bbu(t)}_{L^2}^2
        +2\int_0^t \sen{\n\bbu(s)}_{L^2}^2\rd s
        \leq \sen{\bbu_0}_{L^2},\quad\mathrm{a.e.}\quad t\in [0,T].
        \eex
    \en
    \end{defn}

    Now, our main result reads:
    \begin{thm}\label{thm:main}
    Let $\bbu_0\in L^2(\bbR^3)$ satisfying $\n\cdot\bbu_0=0$,
    $T>0$. Assume that $\bbu$ is a weak solution of \eqref{NSE} in $[0,T]$.
    If there exists an absolute constant $\ve_0>0$ such that
    \be\label{thm:main:reg}
    \sen{\bbu}_{\dot B^{-1}_{\infty,\infty}}\leq\ve_0,
    \ee
    then $\bbu$ is smooth in $(0,T)$.
    \end{thm}

    The rest of this paper is organized as follows. In section \ref{sect:Pre},
    we recall the definition of
    Besov spaces and an interpolation inequality.
    Section \ref{sect:proof} is devoted to proving  Theorem \ref{thm:main}.

\section{Preliminaries}\label{sect:Pre}

    We first introduce the Littlewood-Paley decomposition. Let $\calS(\bbR^3)$ be the Schwartz class of rapidly decreasing functions. For $f\in \calS(\bbR^3)$, its Fourier transform $\calF f=\hat f$ is defined as
    \bex
    \hat f(\xi)=\int_{\bbR^3}f(x)e^{-ix\cdot \xi}\rd x.
    \eex

    Let us choose an non-negative radial function $\varphi\in \calS(\bbR^3)$ such that
    \bex
    0\leq \hat \varphi(\xi)\leq 1,\quad \hat \varphi(\xi)=\left\{\ba{ll}
    1,&\mbox{if }|\xi|\leq 1,\\
    0,&\mbox{if }|\xi|\geq 2,
    \ea\right.
    \eex
    and let
    \bex
    \psi(x)=\varphi(x)-2^{-3}\varphi(x/2),\
    \varphi_j(x)=2^{3j}\varphi(2^jx),\
    \psi_j(x)=2^{3j}\psi(2^jx),\quad j\in\bbZ.
    \eex
    For $j\in\bbZ$, the Littlewood-Paley projection operators $S_j$ and $\lap_j$ are, respectively, defined by
    \bex
    S_jf=\varphi_j*f,\quad \lap_jf=\psi_j*f.
    \eex
    Observe that $\lap_j=S_j-S_{j-1}$. Also, it is easy to check that if $f\in L^2(\bbR^3)$, then
    \bex
    S_jf\to 0,\mbox{ as }j\to -\infty;\quad
    S_jf\to f,\mbox{ as }j\to \infty,
    \eex
    in the $L^2$ sense. By telescoping the series, we have the following Littlewood-Paley decomposition
    \bex
    f=\sum_{j=-\infty}^\infty\lap_jf,
    \eex
    for all $f\in L^2(\bbR^3)$, where the summation is in the $L^2$ sense.

    Let $s\in\bbR$; $p,q \in [1,\infty]$, the homogeneous Besov space $\dot B^s_{p,q}(\bbR^3)$ is defined by the full dyadic decomposition such as
    \bex
    \dot B^s_{p,q}
    =\sed{f\in \calZ'(\bbR^3);\ \sen{f}_{\dot B^s_{p,q}}=\sen{\sed{
    2^{js}\sen{\lap_jf}_{L^p}
    }_{j=-\infty}^\infty}_{\ell^q}<\infty},
    \eex
    where $\calZ'(\bbR^3)$ is the dual space of
    \bex
    \calZ(\bbR^3)=\sed{f\in\calS(\bbR^3);\
    D^\alpha \hat f(0)=0,\quad \forall\ \alpha\in \bbN^3}.
    \eex

    The following interpolatin inequality will be need in Section \ref{sect:proof},
    \bee\label{inter_ineq}
    \sen{f}_{L^q}
    \leq C
    \sen{f}_{\dot H^{\alpha\sex{\frac{q}{2}-1}}}^\frac{2}{q}
    \sen{f}_{\dot B^{-\alpha}_{\infty,\infty}}^{1-\frac{2}{q}},\quad
    \forall \ f\in \dot H^{\alpha\sex{\frac{q}{2}-1}}(\bbR^3)
    \cap \dot B^{-\alpha}_{\infty,\infty}(\bbR^3),
    \eee
    where $2<q<\infty$ and $\alpha>0$. See \cite{MGO_97} for the proof.

\section{Proof of Theorem \ref{thm:main}}\label{sect:proof}

    In this section, we shall prove Theorem \ref{thm:main}.

    By the classical ``weak=strong" type uniqueness theorem, we need only to derive the a priori estimate
    \bee\label{goal}
    \bbu\in L^\infty(0,T;H^1(\bbR^3))\cap L^2(0,T;H^2(\bbR^3)).
    \eee

    Multiplying $\eqref{NSE}_1$ by $-\lap \bbu$, integrating over $\bbR^3$, we obtain
    \bee\label{energy}
    \bea
    \frac{1}{2}
    \frac{\rd}{\rd t} \sen{\n\bbu}_{L^2}^2
    +\sen{\lap\bbu}_{L^2}^2
    &=\int_{\bbR^3} [(\bbu\cdot\n)\bbu]\cdot\lap\bbu\rd x\\
    &\equiv I.
    \eea
    \eee
    By H\"older inequality,
    \bex
    I\leq \sen{\bbu}_{L^6}\sen{\n\bbu}_{L^3}
        \sen{\lap\bbu}_{L^2}.
    \eex
    Invoking \eqref{inter_ineq} with $q=6$, $\alpha=1$; and $q=3$, $\alpha=2$, we may further estimate $I$ as
    \bee\label{I}
    \bea
    I&\leq C\sex{\sen{\bbu}_{\dot H^2}^\frac{1}{3}
        \sen{\bbu}_{\dot B^{-1}_{\infty,\infty}}^\frac{2}{3}}
        \sex{\sen{\n\bbu}_{\dot H^1}^\frac{2}{3}\sen{\n\bbu}_{\dot B^{-2}_{\infty,\infty}}^\frac{1}{3}}
        \sen{\lap\bbu}_{L^2}\\\
    &=C\sen{\bbu}_{\dot B^{-1}_{\infty,\infty}}\sen{\lap\bbu}_{L^2}^2.
    \eea
    \eee
    Substituting \eqref{I} into \eqref{energy}, we see
    \bex
    \frac{1}{2}
    \frac{\rd}{\rd t} \sen{\n\bbu}_{L^2}^2
    +\sex{1-C\sen{\bbu}_{\dot B^{-1}_{\infty,\infty}}}\sen{\lap\bbu}_{L^2}^2
    \leq 0.
    \eex
    Thus, if
    \bex
    \sen{\bbu}_{\dot B^{-1}_{\infty,\infty}}\leq \frac{1}{C}\equiv \ve_0,
    \eex
    we deduce that $\sen{\n\bbu}_{L^2}$ is decreasing, and thus \eqref{goal}, as desired.

    The proof of Theorem \ref{thm:main} is completed.

\end{document}